OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4) i kind of remember but im not sure

OpenStudy (anonymous):

m=2

OpenStudy (anonymous):

\[\frac{ 4-2 }{ 7-6 }=\frac{ 2 }{ 1 }\]

OpenStudy (anonymous):

so yes, m=2

OpenStudy (anonymous):

thank youu